# Probing Chiral-Symmetric Higher-Order Topological Insulators with Multipole Winding Number

### 08 Jan 2024 v1

@@ Abstract
Figure 3: MWN obtained from the CTBC [Eq. (1)] (black dots), the non-commutative form [Eq. (8)] (blue dots), and the time-evolution [Eq. (11)] (red circles). Bulk initial states are chosen as Fock states ranging from $[L/3,2L/3]\times[L/3,2L/3]$. For initial states on the edge, we find it is sufficient to use four Fock states localized at the center of the edge: $(1,L/2),(L/2,1),(L,L/2),(L/2,L)$. Insets: the time-evolution of $\langle\Psi(t)|\hat{\Gamma}\hat{\cal P}_{xy}|\Psi(t)\rangle$ with the red dashed lines for the edge state and the blue lines for the bulk state. The system size is chosen as $30\times 30$, and the total evolution time is set as $\gamma t=3$.
In practice, one may use dynamical average to obtain the expectation over eigenstates. Below we show how to measure the MWN for a 2D BBH model. Under the CTBC, it can be found that there are boundary states localized near the four sides with non-zero energy. Hence, the MWN should not only relate to the bulk states, but also the boundary states. To attain Eq. (8), contributions from bulk and boundary states should be both taken into account. The numbers of boundary and bulk states are roughly $N_{\mathrm{edge}}=4L$ and $N_{\mathrm{bulk}}=L^{2}-4L$. We consider two kinds of initial states: the edge initial state $|{\Psi_{\mathrm{edge}}}\left(0\right)\rangle=\sum\nolimits_{\psi_{n}\in\mathrm{edge}}{{b_{n}}|{\psi_{n}}\rangle}$ and the bulk initial state $|{\Psi_{\mathrm{bulk}}}\left(0\right)\rangle=\sum\nolimits_{\psi_{n}\in\mathrm{bulk}}{{a_{n}}|{\psi_{n}}\rangle}$. Based upon the time-evolution from these initial states, one can obtain the contribution from the bulk
@@
Through numerical simulation, we check the validity of Eq. (11). We simulate the time-evolution from Fock states that localize at the center of the bulk and the edge, respectively. Then, we calculate $\nu_{\mathrm{bulk}}$ and $\nu_{\mathrm{edge}}$ to obtain the MWN via Eq. (11). These procedures are repeated for different Fock states chosen appropriately to ensure that all eigenstates except for the zero-energy modes are covered. By averaging them, we obtain the results in Fig. 3. The simulated results are consistent with the ones given by the non-commutative form Eq. (8). Note that it is difficult to achieve a very long time evolution in realistic experiments. In our scheme, it is sufficient to consider the evolution before states reach the corners, which avoids the influence of corner states (see details in Supplemental Material [33]).
@@
Summary and Discussion. In summary, to characterize chiral-symmetric HOTIs, we have introduced the MWN and explored its experimental measurement. Through introducing suitable twist operators to transform the gauge field at corners into the bulk, the twist angle serves as a perturbation, leading to the derivation of a real-space formula for the MWN. This formula provides a practical approach for experimentally probing higher-order topology. Utilizing the corner twisted boundary condition, we demonstrate the effectiveness of the TBC-like method in computing real-space topological invariants.
@@
In contrast to the conventional TBC, the CTBC exclusively connects corners, leading to the emergence of boundary states near borders in non-trivial cases. Notably, these boundary states have non-zero energies. The MWN is rooted in both bulk and edge contributions, setting it apart fundamentally from conventional chiral-symmetric topological insulators. This distinction necessitates measuring both contributions in experiments. Furthermore, the appearance of corner states is a consequence of the non-trivial topology of both the bulk and the edge. Thus, unlike the conventional topological band theory, the MWN can not be solely characterized with the Bloch states in bulk under the PBC.
@@
In our real-space formulation, the twist operators [Eq. (2) and Eq. (Probing Chiral-Symmetric Higher-Order Topological Insulators with Multipole Winding Number)] significantly differ from the commonly used quadrupole and octupole [36, 37, 38, 24, 25]. This distinction arises from their connection to the topological invariant defined through specific TBC (see details in Supplemental Material [33]). Due to the gauge freedom, the twist operator is not unique, allowing for numerous gauge transformations and resulting in various possible real-space formulations using the introduced perturbative expansions. Despite these differences, they all share the same physical origin, underscoring the importance of the TBC method in studying real-space formulas for topological invariants. Moreover, the MWN defined through the CTBC effectively captures information about zero-energy modes. This enables the generalization of our method to a broad range of chiral-symmetric HOTIs beyond square or cubic structures, such as the honeycomb lattice [39, 40, 41, 42] and the kagome lattice [8, 43, 7]. The real-space formula for the corresponding MWN can be derived in a similar manner.
Acknowledgements.

### 21 Mar 2024 v2

@@ Results
where $\hat{\mathcal{P}}=\{\hat{\cal{P}}_{xy},\hat{\cal{P}}_{xyz}\}$.
We note that there are boundary states localized near each side under both CTBC and OBC. As mentioned above, these states possess non-zero energy and, consequently, contribute to the computation of MWN. To elucidate this phenomenon, we evaluate $\langle\psi_{n}|\hat{\Gamma}\hat{\mathcal{P}}_{xy}|\psi_{n}\rangle$ for each eigenstate in 2D BBH model, as depicted in Fig. 4. Notably, boundary states and bulk states exhibit distinct contributions, manifesting the feature of higher-order topology compared to the conventional topological insulator.
Figure 4: Contribution of each eigenstates to the MWN in 2D BBH model. a The expectation of the chiral multipole operator $\langle\psi_{n}|\hat{\Gamma}\hat{\mathcal{P}}_{xy}|\psi_{n}\rangle$ for each eigenstate except for zero-energy modes in 2D BBH model. Red circles and blued dots are distinguished by its value. The indices of eigenstates are arranged in ascending order of energy. b-c respectively shows the total density distribution of eigenstates marked by red and blue colors in (a). Parameters are chosen as $\gamma/\lambda=0.5$ and $L=30$.
Measurement of multipole winding number through dynamical evolution. In practice, one may use dynamical average to obtain the expectation over eigenstates. Below we show how to measure the MWN for a 2D BBH model. Under the CTBC, it can be found that there are boundary states localized near the four sides with non-zero energy. Hence, the MWN should not only relate to the bulk states, but also the boundary states. To attain Eq. (14), it is crucial that the contributions from bulk and boundary states should be both taken into account. In the BBH model, the numbers of boundary and bulk states are roughly $N_{\mathrm{edge}}=4L$ and $N_{\mathrm{bulk}}=L^{2}-4L$. We consider two kinds of initial states: the edge initial state $|{\Psi_{\mathrm{edge}}}\left(0\right)\rangle=\sum\nolimits_{\psi_{n}\in\mathrm{edge}}{{b_{n}}|{\psi_{n}}\rangle}$ and the bulk initial state $|{\Psi_{\mathrm{bulk}}}\left(0\right)\rangle=\sum\nolimits_{\psi_{n}\in\mathrm{bulk}}{{a_{n}}|{\psi_{n}}\rangle}$. Based upon the time-evolution from these initial states, one can obtain the contribution from the bulk
@@
Through numerical simulation, we check the validity of Eq. (12). We simulate the time-evolution from Fock states that localize at the center of the bulk and the edge, respectively. Then, we calculate $\nu_{\mathrm{bulk}}$ and $\nu_{\mathrm{edge}}$ to obtain the MWN via Eq. (12). These procedures are repeated for different Fock states chosen appropriately to ensure that all eigenstates except for the zero-energy modes are covered. By averaging them, we obtain the results in Fig. 5. The simulated results are consistent with the ones given by the non-commutative form Eq. (14). Note that it is difficult to achieve a very long time evolution in realistic experiments. In our scheme, it is sufficient to consider the evolution before states reach the corners, which avoids the influence of corner states (see discussions in Sec. D in Supplementary Note 2).
Figure 5: Numerical simulation of measuring MWN from dynamical evolution. MWN obtained from the CTBC [Eq. (1)] (black dots), the non-commutative form [Eq. (14)] (blue dots), and the time-evolution [Eq. (12)] (red circles). Bulk initial states are chosen as Fock states ranging from $[L/3,2L/3]\times[L/3,2L/3]$. For initial states on the edge, we find it is sufficient to use four Fock states localized at the center of the edge: $(1,L/2),(L/2,1),(L,L/2),(L/2,L)$. Insets: the time-evolution of $\langle\Psi(t)|\hat{\Gamma}\hat{\cal P}_{xy}|\Psi(t)\rangle$ with the red dashed lines for the edge state and the blue lines for the bulk state. The system size is chosen as $30\times 30$, and the total evolution time is set as $\gamma t=3$.
Discussions
@@
In summary, to characterize chiral-symmetric HOTIs, we have introduced the MWN and explored its experimental measurement. Through introducing suitable twist operators to transform the gauge field at corners into the bulk, the twist angle serves as a perturbation, leading to the derivation of a real-space formula for the MWN. This formula provides a practical approach for experimentally probing higher-order topology. Particularly, our method may be applied to those chiral-symmetric HOTIs with higher winding number in the presence of long-range tunneling [25, 28, 29], in which the number of corner states can be hardly distinguished in large systems. By utilizing the corner twisted boundary condition, we demonstrate the effectiveness of the TBC-like method in studying real-space topological invariants and the bulk-corner correspondence.
@@
In contrast to the conventional TBC, the CTBC exclusively connects corners, leading to the emergence of boundary states near borders in non-trivial cases. Notably, these boundary states have non-zero energies. The MWN is rooted in both bulk and boundary states, setting it apart fundamentally from conventional chiral-symmetric topological insulators. This distinction necessitates measuring both contributions from the bulk and boundary in experiments. Furthermore, the appearance of corner states is a consequence of the non-trivial topology of both the bulk and the edge. Thus, unlike the conventional topological band theory, the MWN can not be solely characterized with the Bloch states in bulk under the PBC. This also explains why the MWN can well identify different BOTPs.
@@
In our real-space formulation, the twist operators [Eq. (3) and Eq. (Results)] significantly differ from the commonly used quadrupole and octupole operators [40, 41, 42, 23, 24]. This distinction arises from their connection to the topological invariant defined through specific TBC (see detailed derviations in Sec. A in Supplementary Note 2). Due to the gauge freedom, the twist operator is not unique, allowing for numerous gauge transformations and resulting in various possible real-space formulations using the introduced perturbative expansions. Despite these differences, they all share the same physical origin, underscoring the importance of the TBC method in studying real-space formulas for topological invariants. Moreover, the MWN defined through the CTBC effectively captures information about zero-energy modes. This enables the generalization of our method to a broad range of chiral-symmetric HOTIs beyond square or cubic structures, such as the honeycomb lattice [43, 44, 45, 46], the kagome lattice [9, 47, 48, 49, 50] and even the quasi crystalline lattice [51, 52, 53, 54]. On the other hand, it is also potential to generalize the MWN to the Floquet HOTIs [39, 38, 55, 56, 57, 58] with appropriate construction. The real-space formula for the corresponding MWN can be derived in a similar manner.
Methods
Chiral-symmetric lattice systems and singular value decomposition. Usually, a chiral-symmetric lattice system can be divided into two sublattices obeying a tight-binding Hamiltonian:
$\hat{H}={{\bm{\psi}}_{A}}h{\bm{\psi}}_{B}^{\dagger}+{{\bm{\psi}}_{B}}{h^{\dagger}}{\bm{\psi}}_{A}^{\dagger},$ (13)
where ${{\bm{\psi}}_{\alpha}}\equiv\left({|{1_{\alpha}}\rangle,|{2_{\alpha}}\rangle,\cdots}\right)$ with $\alpha=\{A,B\}$ indexing the two sublattices and $|i_{\alpha}\rangle$ being the single-particle position-space basis for the $\alpha$ sublattice. For convenience, we assume the dimensions of $A$ and $B$ sublattices are equal. The Hamiltonian is completely determined by the off-diagonal matrix $h$. The chiral symmetry results in a symmetric spectrum: the eigenvalues of $H$ appear pairwise with opposite values, which are related to the singular values of $h$ [27]. The eigen-equation $\hat{H}|{\Psi_{n}}\rangle={E_{n}}|{\Psi_{n}}\rangle$ can be expressed as two coupled equations: $hu_{B}^{[n]}={E_{n}}u_{A}^{[n]}$ and ${h^{\dagger}}u_{A}^{[n]}={E_{n}}u_{B}^{[n]}$. Here, the $n$th eigenstate reads as $|{\Psi_{n}}\rangle=\frac{1}{{\sqrt{2}}}\sum\nolimits_{j}{({u_{A}^{j,n}|{j_{A}}\rangle+u_{B}^{j,n}|{j_{B}}\rangle})}$, and the column vectors $u_{A,B}^{[n]}$ have $u^{j,n}_{A,B}$ as their $j$th elements. By performing the singular value decomposition: $h={u_{A}}\Sigma{u_{B}}^{-1}$, where $\Sigma=\mathrm{diag}(E_{1},\cdots,E_{n},\cdots)$ is diagonal-positive-definite and $u_{A,B}=[u_{A,B}^{[1]},u_{A,B}^{[2]},\cdots,u_{A,B}^{[L/2]}]$ is an unitary matrix, one can solve the coupled equations. It can be seen that the singular values of matrix $h$ are exactly half of the energy spectrum.
Derivation of the real-space formula of multipole winding number. Choosing $r=1$, under the bulk gauge, the MWN can still be expressed as Eq. (1). Keeping the terms up to the first order of $1/L$, we have $\tilde{h}\left(\theta\right)=h\left(0\right)+\frac{\theta}{L}{[{\frac{\partial}{{\partial\theta/L}}\tilde{h}\left(\theta\right)}]_{\theta=0}}+O\left({\frac{1}{{{L^{2}}}}}\right)$ and ${\rm{Tr}}[{h{{\left({r,\theta}\right)}^{-1}}{\partial_{\theta}}h\left({r,\theta}\right)}]\approx{\rm{Tr}}\{{{h^{-1}}\left(0\right){{[{{{\partial_{\theta}}}\tilde{h}\left(\theta\right)}]}_{\theta=0}}}\}$ being a constant. Therefore, the MWN can be given as $\nu\approx-i{\rm{Tr}}\{{{h^{-1}}\left(0\right){{[{{{\partial_{\theta}}}\tilde{h}\left(\theta\right)}]}_{\theta=0}}}\}$. Using $\tilde{h}\left(\theta\right)={\bm{\psi}}_{A}^{\dagger}\tilde{H}\left(\theta\right){{\bm{\psi}}_{B}}$, we have $\tilde{h}={\cal U}_{2\pi}^{A}\tilde{h}\left(0\right){({{\cal U}_{2\pi}^{B}})^{-1}}$ with ${{\hat{U}}_{2\pi}}=\exp({i2\pi\hat{\mathcal{P}}/L})$, where $\hat{\mathcal{P}}=\{\hat{\cal{P}}_{xy},\hat{\cal{P}}_{xyz}\}$ are both diagonal in the position basis. For simplicity, we define the projected twist operator matrices in the sector $\alpha=\{A,B\}$: ${\cal U}_{2\pi}^{\alpha}\equiv{\bm{\psi}}_{\alpha}^{\dagger}{{\hat{U}}_{2\pi}}{{\bm{\psi}}_{\alpha}}$. To ensure the gauge invariance, we impose the constraint $\det\tilde{h}\left({2\pi}\right)=\det\tilde{h}\left(0\right)$. Thus, we require ${\cal U}_{2\pi}\equiv{\cal U}_{2\pi}^{A}={\cal U}_{2\pi}^{B}$, which can be easily satisfied by setting an identical position for all sublattices within a unit cell. Up to the first order of $1/L$, we have ${\left[{{\partial_{\theta}}\tilde{h}\left(\theta\right)}\right]_{\theta=0}}\approx\frac{{\tilde{h}\left({2\pi}\right)-\tilde{h}\left(0\right)}}{{2\pi}}=\frac{{{\cal U}_{2\pi}h{{{{\cal U}_{2\pi}^{-1}}}}-h}}{{2\pi}}$. To gain higher accuracy, one may consider a higher-order finite difference (see details in Sec. B of Supplementary Note 2). Then, a rather simpler real-space formula can be obtained: $\nu=\frac{1}{{2\pi i}}{\rm{Tr}}\left({{h^{-1}}{\cal U}_{2\pi}h{{{{\cal U}_{2\pi}^{-1}}}}-I}\right)$. This leads to two equivalent real-space formulas: (i) the Bott index form and (ii) the non-commutative form.
Since ${{h^{-1}}{\cal U}_{2\pi}h{{{{\cal U}_{2\pi}^{-1}}}}}$ is close to an identity matrix, in the thermodynamic limit $L\to\infty$, one can obtain the Bott index form $\nu=\frac{1}{{2\pi i}}{\rm{Tr}}\left[{\log\left({{h^{-1}}{\cal U}_{2\pi}h{{{{\cal U}_{2\pi}^{-1}}}}}\right)}\right]$. This formula is analogous to the multipole chiral number [22], which also takes the Bott index form. The multipole chiral number is defined based on the quadrupole and octupole of polarization. It can be proved that the multipole chiral number is equivalent to the real-space winding number defined through a modified CTBC (see details in Sec. A of Supplementary Note 2).. The modified CTBC still connects the corners, while only those tunnelings related to one of the corners are threaded by a twist angle. In this scenario, the modified MWN captures the number of zero-energy modes induced in one specific corner. Therefore, although the generators of the twist operators [Eq. (3) and Eq. (Results)] are quite different from the quadrupole and octupole, they share the same physical origin.
Using the Baker-Campbell-Hausdorff (BCH) formula, we have ${{\cal U}_{2\pi}}h{{\cal U}_{2\pi}^{-1}}\approx h+2\pi i\left[{P,h}\right]/L$ with $P={\bm{\psi}}_{A}^{\dagger}\hat{\mathcal{P}}{\bm{\psi}}_{A}$ denoting the generator projected onto the sector $A$, and then obtain the non-commutative form $\nu=\frac{1}{L}\mathrm{Tr}\left({h^{-1}}\left[{{{P}},h}\right]\right)$. However, this approximation only holds for systems with OBC and $L\to\infty$. In a finite system with CTBC, this approximation is poor because it always evaluates to zero if $h$ is invertible: ${\rm{Tr}}\left({{h^{-1}}\left[{P,h}\right]}\right)={\rm{Tr}}\left({{h^{-1}}Ph-P}\right)=0$. This issue arises from the slow convergence of the BCH formula and the omission of higher-order terms (see discussions in Sec. B of Supplementary Note 2). Meanwhile, as the CTBC connects the corners, general definition of the position near the corner become ill-defined, which also leads to the poor performance of the first-order expansion. However, under the OBC, the BCH formula converges fast because position-dependent operators are well-defined. In the presence of zero modes, $h$ becomes singular. In this case, we need to rule out the zero singular values and the associated vectors and use the pseudo-inverse formulation: $h^{-1}\approx\tilde{u}_{B}\tilde{\Sigma}^{-1}\tilde{u}_{A}^{{\dagger}}$ and then the non-commutative form will be non-zero.
The above forms for MWN are not straightforward to be measured in experiments. Below we derive a formula suitable for experimental measurement. Firstly, the flattened Hamiltonian can be written as $\hat{Q}={{\hat{\mathbb{P}}}_{+}}-{{\hat{\mathbb{P}}}_{-}}$ with ^${{\hat{\mathbb{P}}}_{-}}=\sum\nolimits_{{E_{n}}<0}{|{\psi_{n}}\rangle\langle{\psi_{n}}|}$ and ^${{\hat{\mathbb{P}}}_{+}}=\sum\nolimits_{{E_{n}}>0}{|{\psi_{n}}\rangle\langle{\psi_{n}}|}$, where $E_{n}$ is the eigenenergy for the eigenstate $|\psi_{n}\rangle$. The flattened Hamiltonian shares the same eigenstates of the original Hamiltonian and the zero-energy mode is naturally excluded. Replacing the off-diagonal matrix $h$ with the flattened matrix $q\equiv{\bm{\psi}}_{A}^{\dagger}\hat{Q}{{\bm{\psi}}_{B}}$ and $q^{-1}\equiv{\bm{\psi}}_{B}^{\dagger}\hat{Q}{{\bm{\psi}}_{A}}$, we have $\nu=\frac{1}{L}\mathrm{Tr}\left({q^{-1}}\left[{{{P}},q}\right]\right)$. Therefore, we obtain a real-space formula (see detailed derivations Sec. C of Supplementary Note 2),
Acknowledgments